According to the Cable News Network (CNN), the number of injured in-line skaters (or “roller-bladers”) was 184% larger in 1994 than it was in 1993. Did the number of injured skaters almost double, almost triple, or more than triple?
Answer by edjones(7569)

Let the initial number=100
100 * 1.84 = 184
100+184=284
284/100=2.84 almost triple.
.
Ed